If alpha squared =5*alpha-3, beta squared =5*beta-3 then value of alpha/beta+beta/alpha is?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,

As Given,
  alpha + beta = 5
       ( Alpha)*(beta)= 3
  •         = { (alpha+beta)^2 – 2alpha*beta } / alpha*beta
 =( 25-6 ) / 3
 = 19/3
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)